How much power is needed to lift a 4.5 kg object to a height of 3.0 m in 12 s?

1
Expert's answer
2022-03-07T07:16:07-0500

Power

"P=\frac{W}{t}=\frac{mgh}{t}"

"P=\frac{4.5\times9.8\times3}{12}=11.025W"